OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (paxpolaris):

first find the slope:

OpenStudy (paxpolaris):

\[\Large m={1-2 \over 2-(-3)}= -\frac15\]

OpenStudy (paxpolaris):

if you know the slope(m), and the line passes through a point (h,k) then the equation is: y-k=m(x-h)

OpenStudy (paxpolaris):

so you get: \[\LARGE y-2= -\frac15\left(x+3\right)\]or\[ \LARGE y-1= -\frac15\left(x-2\right)\]

OpenStudy (anonymous):

thank you